# Learning on LLM Output Signatures for gray-box LLM Behavior Analysis

### v1

@@ 5 Experiments
We assess various aspects of learning with LOS via the following questions: (Q1) Is learning on LOS an effective approach for addressing key tasks such as DCD and HD? Does it outperform baselines? (Sections 5.2 and 5.1); (Q2) Does our model exhibit transfer capabilities across LLMs and across datasets, suggesting the emergence of universal patterns in LLM behavior from the LOS perspective? (Section 5.3); (Q3) How important is $\mathbf{X}$ in the pair $(\mathbf{X},\mathbf{p})$, as it is often overlooked? And how impactful is the choice of the slicing parameter $K$ in Equation 1 (Section C.6). In the following, we present our main results, and refer to Appendix C for additional experiments, and experimental details.
@@
Our experiments focus on the two tasks of DCD and HD, with hyperparameter $K$ fixed at $1000$ unless stated otherwise (see Equation 1). In Section C.6, we show that our model is robust to variations in $K$. To align with prior work, we use datasets and LLMs from [41, 56] for DCD and [35] for HD, totaling six datasets – three for DCD and three for HD – and seven LLMs, five for DCD and two for HD. Further details are in subsequent sections. We use the area under the ROC curve (AUC) to evaluate HD and DCD, a standard metric in this domain [35, 41, 56], which measures the balance between sensitivity and specificity. We conduct each experiment across three different random seeds (when applicable) and report the mean along with the standard deviation of the results.
Newly introduced learning-based baselines. In addition to task-specific baselines, detailed in the following, we also introduce two novel learning-based baselines to appreciate the contribution of TDS. In these baselines, which we call ATP+Rank-MLP and ATP+Rank-Transformer (dubbed ATP+R-MLP, ATP+R-Transf., respectively), we ablate information about the TDS: they only process Rank Encodings (Equation 2), thus accessing the ATP and rank information only. Formal definitions are in Section C.4.
Table 1: Comparison of AUC over Mis-7b and L3-8b on HD, across the discussed baseline methods. The best-performing method is in bold, and the second best is underlined.
@@
96.11
$\pm$ 0.03
68.59
$\pm$ 1.08
72.97
$\pm$ 0.41
89.44
$\pm$ 0.32
77.04
$\pm$ 0.77
Table 2: Test AUC on BookMIA. ‘P’: Pythia, ‘L’: LLaMa-1. Best result is in Bold, second best is underlined. Reference-based approaches are shaded in pink.
@@
5.1 Hallucination Detection
We follow the setup of Orgad et al. [35]. The objective is to predict whether an LLM-generated response to a given input prompt is correct or not. We frame the task within a gray-box setting, i.e., we assume no access to the LLM’s internals. We also assume no access to external resources, e.g., other LLMs, auxiliary, repeated prompting, or any additional contextual information, such as pointers to specific answer tokens. This makes methods like [35, 23] not directly comparable. Note, however, that these additional sources of information could easily be incorporated into LOS-Net, e.g., using a one-hot vector to flag specific answer tokens or by extending the LOS to account for additional prompting. We defer investigating these more relaxed settings to future research.
Datasets and LLMs. Following Orgad et al. [35], we use three datasets spanning various domains and tasks: HotpotQA without context [52], IMDB sentiment analysis [30], and movie roles [35]. Further details, regarding annotations’ collection process, the splits and dataset sizes are in Section C.5.1. As the target LLMs, coherently with Orgad et al. [35], we use Mistral-7b-instruct-v0.2 [21] (Mis-7b), and LLaMa3-8b-instruct [46] (L-3-8b).
HD Baselines. As baselines for hallucination detection, we consider the following, (1) Aggregated probabilities/logits: Previous studies [14, 22, 48, 18] simply aggregate output token probabilities or logits to score LLM confidence for error detection. Such simple aggregations include mean/max/min over the ATP. We refer to them as Logit/Probas-mean/min/max; (2) P(True): [22] found that LLMs show reasonable calibration in assessing their own output correctness.
Results. Table 1 presents a comprehensive summary of our main results, which clearly demonstrate that LOS-Net outperforms all baselines across all six dataset/LLM combinations, often by a significant margin. For instance, on the IMDB dataset, LOS-Net achieves an AUC improvement of around 34 units over the best baseline for Mis-7b and 16 over the best baseline for L-3-8b-instruct. Our results further indicate that ATP learning-based baselines consistently underperform compared to LOS-Net, underscoring the critical role of the TDS, $\mathbf{X}$, in achieving superior results. However, our ATP-based learnable baselines outperform Probas/Logits-based baselines in 3 out of 6 cases, suggesting that a learning approach relying exclusively on ATP can still be a viable solution in certain scenarios.
Dataset $\rightarrow$ WikiMIA - 32 WikiMIA - 64
LLM $\rightarrow$ P-6.9b L-13b L-30b M-1.4b P-6.9b L-13b L-30b M-1.4b
5.2 Data Contamination Detection
Datasets and LLMs. We use three datasets to assess DCD, specifically: WikiMIA-32 and WikiMIA-64 [41], as well as BookMIA [41]. The WikiMIA-32 and -64 datasets contain excerpts from Wikipedia articles, consisting of, resp., 32 and 64 words. The distinction between contaminated and uncontaminated data is determined by timestamps. As in [41, 56], we attack Mamba-1.4b [13] (M-1.4b), LLaMa-13b/30b [46] (L-13b/30b), Pythia-6.9b [4] (P-6.9b). BookMIA is a dataset of book excerpts. Positive members correspond to books known to be well memorized by certain OpenAI models [11], or otherwise known to (partly) be in pretraining corpus of other open-source LLMs [1]. Non-members include excerpts from books released after 2023, necessarily absent from the pretraining corpus of the these last ones. Interestingly, this dataset allows us to test LOS-Net’s DCD capability in a realistic scenario akin to copyright-infringement detection. We thus propose a new split that ensures all excerpts from the same book always appear either in the training or test split (and never in both). Details are enclosed in Section C.5.2. We attack LLMs considered in [1]: LLaMa-13b/30b [46] (L-13b/30b), Pythia-6.9b/12b[4] (P-6.9b/12b).
DCD Baselines. We evaluate six recent methods as our baselines. The Loss approach [53] directly uses the loss value as the detection score. The Reference (Ref) method [9] calibrates the target LLM’s perplexity leveraging a similar reference model known or supposed not to have memorized text of interest [^4] For example for Pythia-12b, a valid reference LLM would be the smaller Pythia-70M.. Both Zlib and Lowercase [9] are also reference-based methods: they utilize zlib compression entropy and lowercased text perplexity as reference for normalization. Lastly, Min-K[41] and Min-K%++ [56] are reference-free methods, which examine token probabilities and average a subset of the minimum token scores, or a function thereof, over the input. Min-K%++ is currently the best-performing method on the WikiMIA dataset. For these baselines, we select their hyperparameters by maximizing performance on the validation set(s).
@@ 5.2 Data Contamination Detection
Results are reported in Table 2. On this benchmark, our method attains exceptional results, largely surpassing other reference-free approaches. Among these last ones, ours is the only method that can match or outperform the reference-LLM-based baselines. Additionally, our experimental results suggest that instrumental to achieve such strong reference-free performance is to access, even partially, the TDS, as our ATP-based learnable baselines, which only process features for the actual sequence tokens, incur significant performance degradations.
[…]
As shown in Table 3, LOS-Net consistently surpasses all baseline methods across all eight combinations of LLMs and datasets. Notably, for L-30b, our model achieves an AUC score that is more than 8 points higher than the best-performing baseline, MinK%++ for both datasets, demonstrating a substantial improvement. Similarly, for P-6.9b, our model maintains a steady advantage of approximately 5 AUC for both datasets, further underscoring its robustness. Overall, the second-best method is MinK%++, followed by MinK%, consistently with the findings of [56].
@@ 5.3 Generalization to other LLMs and datasets
Figure 3: BookMIA zero-shot generalization. Bold: highest score among reference-free baselines; superscript : LOS-Net also surpasses reference-based methods.
@@
Here, we further study the possibility to apply our models to settings different from those they were originally trained on. We focus on two variables: datasets and LLMs. Generalization across different datasets was originally studied in [35] within the scope of HD, and in the context of white-box setups; their relevance lies in the fact that non-trivial dataset generalization would potentially suggest a ‘universal truthfulness’ representation encoded in the internal states and/or outputs of an LLM [35, 31, 43]. On the other hand, inspecting transfer across LLMs is, to the best of our knowledge, still unexplored. This study would be important for learning based approaches for applications such as copyright-infringement detection, where ground-truth labels may be scarce.
@@
Zero Shot Cross-LLM Generalization Capabilities in DCD. We assess our model’s ability to detect DC in target LLMs that were unseen during training. Using the BookMIA benchmark and the setup described in Section 5.2, we evaluate our model directly across different LLMs without any fine-tuning. This setup is relevant in cases where contamination information is not yet available for newly released LLMs. The results are presented in the heatmap shown in Figure 3. We observe strong transferability: in 10 out of 12 cases, our model achieves the best performance among reference-free approaches, highlighted in bold in Figure 3. Interestingly, in 3 out of 12 cases, LOS-Net (which is reference-free) even surpasses reference-based baselines, as indicated via a superscript of . We also observe particularly strong transfer across differently sized LLM architectures within the same family and highlight the surprising positive transfer from the largest LLaMa to Pythia models.
Figure 4: Cross-LLM transfer learning performance (AUC). Y-axis: source LLMs (train), X-axis: target LLMs (test). Results better than the best baseline are in Bold; results surpassing training from scratch are marked with .
Figure 5: Cross-dataset transfer learning performance (AUC). Y-axis: train datasets, X-axis: test datasets. Results better than the best baseline are in Bold; results surpassing training from scratch are marked with .
@@
Transfer Learning across LLMs and Datasets for HD. Differently than DCD, where zero-shot application of LOS-Net was successful, for HD we observed non-trivial generalization in the zero-shot setup, however, not sufficient to surpass the simple probability-based techniques. This led us to investigate LOS-Net capabilities in a transfer learning setting, in which we conduct a rapid fine-tuning procedure on all possible LLM/datasets combinations. Specifically, we perform a 10-epoch fine-tuning on the target LLM/dataset (as opposed to 300+ epochs in our standard setting). This process was measured to take less than a minute. We benchmark the fine-tuned model against two baselines. First, to test for successful transfer, we compare with a LOS-Net trained from scratch under an identical setup (i.e., 10 epochs). Second, we contrast the fine-tuned model with the best-reported baseline; this is crucial, as generalization scores above 0.5 AUC are only useful if they surpass non-learnable baselines relying solely on probas/logit outputs.
The test AUC of our fine-tuned LOS-Net’s are reported in Figures 4 and 5. A superscript () indicates that the fine-tuned LOS-Net achieves better performance compared to training from scratch. Bold text highlights cases where the fine-tuned LOS-Net outperforms the best (non-learnable) baseline methods.
@@
Discussion. We observe several key findings. First, LOS-Net exhibits solid transferability in both scenarios. The finetuned models consistently outperform their counterparts trained from scratch: 6/6 cases in the cross-LLM setup (Figure 4), 11/12 cases in the cross-dataset one (Figure 5), as indicated via on the off-diagonal entries; the full results for training from scratch are presented in Figures 8 and 9 in the Appendix. This highlights the effectiveness of LOS as a data type in capturing universal patterns in LLM behavior. Second, from a practical perspective, we find that LOS-Net outperforms the best baseline in 4/6 cases for the cross-LLM scenario (Figure 4) and in 9/12 cases for the cross-dataset scenario (Figure 5), as indicated in bold on the off-diagonal entries. Focusing on the IMDB dataset, when training on L-3-8b and testing on Mis-7b (Figure 4), our model achieves a substantial gain of around 29 AUC units over the best baseline. This result underscores the possibility of transferring across LLMs. A similar trend is observed in the cross-dataset setup (Figure 5): on Mis-7b, when training on HotpotQA or Movies and testing on IMDB, our model achieves a notable improvement of around 31, 27 AUC units, respectively, compared to the best baseline.
6 Conclusion
We proposed LOS-Net, an efficient method for Data Contamination and Hallucinations Detection (DCD, HD) in Large Language Models (LLMs) by leveraging their output signatures (LOS). LOS-Net processes the Token Distribution Sequence (TDS) and Actual Token Probabilities (ATP) using a lightweight transformer with learnable rank encoding, capturing richer contextual information. We theoretically showed that LOS-Net unifies and extends existing gray-box methods under a general framework.
Experiments across datasets and LLMs show LOS-Net outperforms state-of-the-art gray-box baselines in HD and DCD. We demonstrated strong generalization capabilities of LOS-Net, both across datasets and across LLMs, where the latter suggests that LOS-Net can effectively capture universal patterns in LLM behavior. Several avenues for future research remain open. Our framework could extend beyond DCD and HD to other tasks, such as detecting LLM-generated content. Exploring more complex architectures than LOS-Net is also interesting, for example, we note that sorting the TDS tensor removes word alignment across the vocabulary, which may be limiting in some cases.

### v2

@@ 5 Experiments
We assess various aspects of learning with LOS via the following questions: (Q1) Is learning on LOS an effective approach for addressing key tasks such as DCD and HD? Does it outperform baselines? (Sections 5.2 and 5.1); (Q2) Does our model exhibit transfer capabilities across LLMs and datasets, suggesting the emergence of universal patterns in LLM behavior from the perspective of the LOS? (Section 5.3); (Q3) How important is $\mathbf{X}$ in the pair $(\mathbf{X},\mathbf{p})$, as it is often overlooked? And how impactful is the choice of the slicing parameter $K$ in Equation 1? (Section B.6). In the following, we present our main results, and refer to Appendix B for additional experimental results and details.
@@
Our experiments focus on the two tasks of DCD and HD, with hyperparameter $K$ fixed at $1000$ unless stated otherwise (see Equation 1). In Section B.6, we show that our model is robust to variations in $K$. To align with prior work, we use datasets and LLMs from [41, 58] for DCD and [35] for HD, where we also experiment with an additional LLM (Qwen-2.5-7b-Instruct [53]). Further details are in subsequent sections. We use the area under the ROC curve (AUC) to evaluate HD and DCD, a standard metric in this domain [35, 41, 58], which measures the balance between sensitivity and specificity. We conduct each experiment across three different random seeds (when applicable) and report the mean along with the standard deviation of the results. All LOS-Net experiments were conducted using the PyTorch [36] framework on a single NVIDIA L-40 GPU.
Newly introduced learning-based baselines. In addition to task-specific baselines, we also introduce two novel learning-based baselines to appreciate the contribution of the TDS: ATP+R-MLP, ATP+R-Transf.. Specifically, we ablate information about the TDS and only process the ATP and rank information with an MLP or Transformer backbone. Formal definitions are in Section B.4.
5.1 Hallucination Detection
We follow the setup of Orgad et al. [35]. The objective is to predict whether an LLM-generated response to a given input prompt is correct or not. We frame the task within a gray-box setting, i.e., we assume no access to the LLM’s internals. We assume no access to external resources – such as auxiliary LLMs (as in [47]), mulitple queries/generations (as in [35, 23]), or additional contextual cues like token-level annotations ([35]). These approaches exhibit a significantly larger detection latency compared to LOS-Net, problematic in online applications (refer to Section 5.4 for more detailed run-time analyses and comparisons).
Datasets and LLMs. Following Orgad et al. [35], we use three datasets spanning various domains and tasks: HotpotQA without context [54], IMDB sentiment analysis [30], roles in Movies [35]. Details regarding the annotation process, splits and dataset sizes are in Section B.5.1. As the target LLMs, coherently with Orgad et al. [35], we use Mistral-7b-instruct-v0.2 [21] (Mis-7b) and LLaMa3-8b-instruct [46] (L-3-8b), and further experiment with Qwen-2.5-7b-Instruct [53] (Q-2.5-7b).
HD Baselines. (1) Aggregated probabilities/logits: Previous studies [14, 22, 49, 18] simply aggregate output token probabilities or logits to score LLM confidence for error detection. These aggregations operate mean/max/min pooling over the ATP. We refer to them as Logit/Probas-mean/min/max; (2) P(True): [22] found that LLMs show reasonable calibration in assessing their own correctness.
Table 1: Test AUCs for HD over Mis-7b and L3-8b (bold: best method, underlined: second best).
@@ 5.1 Hallucination Detection
77.43
$\pm$ 0.66
Results. Table 1 presents a comprehensive summary of results on the LLMs considered in Orgad et al. [35] (Mis-7b and L-3-8b). These clearly demonstrate that LOS-Net outperforms all baselines across all six dataset/LLM combinations, often by a significant margin. For instance, on the IMDB dataset, LOS-Net achieves an AUC improvement of around 32 units over the best baseline for Mis-7b and 17 over the best baseline for L-3-8b. Our results further indicate that ATP learning-based baselines consistently underperform compared to LOS-Net, underscoring the critical role of the TDS, $\mathbf{X}$. Our ATP-based learnable baselines still outperform non-learnable methods in most cases, suggesting that a learning approach relying exclusively on ATP can still be a viable solution in certain scenarios. Results on Q-2.5-7b are consistent the above findings, and are deferred to Section B.7.
5.2 Data Contamination Detection
Datasets and LLMs. We use three datasets to assess DCD, specifically: WikiMIA-32 and WikiMIA-64 [41], as well as BookMIA [41]. The WikiMIA-32 and -64 datasets contain excerpts from Wikipedia articles, consisting of, resp., 32 and 64 words. The distinction between contaminated and uncontaminated data is determined by timestamps. As in [41, 58], we attack Mamba-1.4b [13] (M-1.4b), LLaMa-13b/30b [46] (L-13b/30b), Pythia-6.9b [4] (P-6.9b). BookMIA is a dataset of book excerpts. Positive members correspond to books known to be well memorized by certain OpenAI models [11], or otherwise known to (partly) be in pretraining corpus of other open-source LLMs [1]. Non-members include excerpts from books released after 2023, necessarily absent from the pretraining corpus of the last ones. Interestingly, this dataset allows us to test LOS-Net’s DCD capability in a realistic scenario akin to copyright-infringement detection. We thus propose a new split that ensures all excerpts from the same book always appear either in the training or test split (and never in both). Details are enclosed in Section B.5.2. We attack LLMs considered in [1]: LLaMa-13b/30b [46] (L-13b/30b), Pythia-6.9b/12b[4] (P-6.9b/12b).
Table 2: Test AUCs on BookMIA. ‘P’: Pythia, ‘L’: LLaMa-1 (bold: best method, underlined: second best, pink: reference-based).
@@ 5.2 Data Contamination Detection
DCD Baselines. The Loss approach [55] directly uses the loss value as the detection score. The Reference (Ref) method [9] calibrates the target LLM’s perplexity leveraging a similar reference model known or supposed not to have memorized text of interest [^4] For example for Pythia-12b, a valid reference LLM would be the smaller Pythia-70M.. Both Zlib and Lowercase [9] are also reference-based methods: they utilize zlib compression entropy and lowercased text perplexity as reference for normalization. Lastly, Min-K[41] and Min-K%++ [58] are reference-free methods, which examine token probabilities and average a subset of the minimum token scores, or a function thereof, over the input. For these baselines, we select their hyperparameters by maximizing performance on the validation set(s).
@@
Refer to Table 2. LOS-Net attains exceptional results, largely surpassing other reference-free approaches. Among these last ones, ours is the only method that can match or outperform even the reference-LLM-based baselines. Importantly, crucial to such strong reference-free performance is to access, even partially, the TDS: our ATP-based learnable methods – which only process features for the actual sequence tokens – incur indeed significant performance drops.
Results on WikiMIA. Due to space limits, the full results on are provided in Section B.8, Table 5. LOS-Net consistently surpasses all baselines across all eight combinations of LLMs and datasets. The second-best method is MinK%++, followed by MinK%, consistent with the findings of [58].
@@ 5.3 Generalization to Other LLMs and Datasets
Figure 3: BookMIA zero-shot generalization (bold: outperforms ref-free baselines, : outperforms ref-based ones).
@@
We further study the possibility to apply our models to settings different from those they were originally trained on. We focus on two variables: datasets and LLMs. Generalization across datasets was originally studied in [35] within the scope of HD and in the context of white-box setups; their relevance lies in the fact that non-trivial dataset generalization would potentially suggest a ‘universal truthfulness’ representation encoded in the internal states and/or outputs of an LLM [35, 31, 43]. Inspecting transfer across LLMs is, to the best of our knowledge, still unexplored. This study is important for learning-based approaches in applications such as copyright-infringement detection, where ground-truth labels may be scarce.
@@
Zero Shot Cross-LLM Generalization Capabilities in DCD. We assess our model’s ability to detect DC in target LLMs that were unseen during training. Using the BookMIA benchmark and the setup described in Section 5.2, we evaluate our model directly across different LLMs without any fine-tuning. This setup is relevant in cases where contamination information is not yet available for newly released LLMs. The results are presented in the heatmap shown in Figure 3. We observe strong transferability: in 10/12 cases, our model achieves the best performance among reference-free approaches, highlighted in bold in Figure 3. Interestingly, in 3/12 cases, LOS-Net (which is reference-free) even surpasses reference-based baselines, as indicated via a superscript of . We also observe particularly strong transfer across differently sized LLM architectures within the same family and highlight the surprising positive transfer from the largest LLaMa to Pythia models.
Figure 4: Cross-LLM transfer Test AUCs (cols: source LLMs, rows: target LLMs). Bold: finetuning LOS-Net outperforms baselines, : it outperforms the same LOS-Net trained from scratch.
Figure 5: Cross-dataset transfer Test AUCs (cols: source data, rows: target data). Bold: finetuning LOS-Net outperforms baselines, : it outperforms the same LOS-Net trained from scratch.
@@
Transfer Learning across LLMs and Datasets for HD. Differently than DCD, where zero-shot application of LOS-Net was successful, for HD we observed non-trivial generalization in the zero-shot setup, however, not sufficient to surpass the simple probability-based techniques. This led us to investigate LOS-Net capabilities in a transfer learning setting, in which we conduct a rapid fine-tuning procedure on all possible LLM/datasets combinations. Specifically, we perform a 10-epoch fine-tuning on the target LLM/dataset (as opposed to 300+ epochs in our standard setting). This process was measured to take less than a minute. We benchmark the fine-tuned model against two baselines. First, to test for successful transfer, we compare with a LOS-Net trained from scratch under an identical setup (i.e., 10 epochs). Second, we contrast the fine-tuned model with the best-reported non-learnable baseline. The test AUC of our fine-tuned LOS-Net’s are in Figures 4 and 5. Superscript ‘’ indicates the fine-tuned LOS-Net is better than a counterpart trained from scratch in the same setting, bold indicates it outperforms the best non-learnable method.
@@
Discussion. First, LOS-Net exhibits solid transferability in both scenarios. The finetuned models consistently outperform their counterparts trained from scratch: 16/18 cases in both the cross-LLM (Figure 4) and cross-dataset setups (Figure 5) – see ‘’ on the off-diagonal entries. This highlights a generally positive transfer of LOS-Net’s learned representations across datasets and LLMs, and underscores the suitability of LOS as a data type in capturing generalizable patterns in LLM behaviors. Second, from a practical perspective, we find that LOS-Net outperforms the best baseline in 15/18 cases for both the cross-LLM (Figure 4) and cross-dataset (Figure 5) scenarios – see bold on the off-diagonal entries. Focusing on the IMDB dataset, when training on L-3-8b and testing on Mis-7b (Figure 4), our model substantially gains around 27 AUC units over the best baseline. This result underscores the possibility of transferring across LLMs. A similar trend is observed in the cross-dataset setup (Figure 5): on Mis-7b, when training on HotpotQA or Movies and testing on IMDB, our model achieves a notable improvement of around 30 AUC units compared to the best baseline.
5.4 Run-Time Analysis
We conclude this section by discussing our comprehensive training and inference timings, reported in Section B.9 and Table 6. We remark how LOS-Net features an extremely contained detection latency: $\approx 10^{-5}$s per inference fwd-pass. Training is also efficient, typically completing in under one hour on a single NVIDIA L-40 GPU, and often taking significantly shorter. To contextualize this computational efficiency w.r.t. methods relying on multiple prompting/generations [35, 23], we measured the generation time of Mistral-7B-Instruct. Averaging over the first 10 samples from the HotpotQA dataset, we observed an average generation time of $1.93\pm 0.18$ seconds per response on a server equipped with eight NVIDIA L-40 GPUs. The methods in [35] and [23] require 5 and 10 generations per detection, resp., rendering their computation overhead around six orders of magnitude higher than that of LOS-Net.
6 Conclusions
We proposed LOS-Net, an efficient method to detect data contamination and hallucinations in LLMs by leveraging their output signatures (LOS), defined as the union of Token Distribution Sequences (TDS) and Actual Token Probabilities (ATP). LOS-Net consists of a lightweight transformer with learnable rank encodings applied on the whole LOS. We proved it unifies and extends existing gray-box methods under a general framework, and experimentally showed it outperforms state-of-the-art gray-box methods across datasets and LLMs. It also exhibited strong generalization capabilities of LOS-Net, both across datasets and across LLMs. Our framework could be applied to other tasks, such as detecting LLM-generated content. Additional sources of information can also be incorporated, e.g., in the absence of latency constraints, it can be interesting to include “exact-token” flags as proposed by [35]. Last, the LOS can be extended to account for multiple prompting [23].
Limitations. By operating in a gray-box setting, our approach is widely applicable, but misses access to predictive information residing, e.g., in the LLM’s hidden states. Additionally, we note that sorting the TDS tensor removes word alignment across the vocabulary, which may be limiting in some cases.